OpenStudy (anonymous):

if a and b are unit vectors and the angle between them is 60degrees calculate |3a-5b| There are unit symbols on a and b. Can someone help me walk through this question?

OpenStudy (anonymous):

Are you familiar with the dot product equation \[a\cdot b=|a||b|\cos \theta?\] Here, a and b are two given vectors, |a| and |b| are the magnitudes/norms of vectors a and b, and theta is the angle between the vectors a and b.